A generic salt, AB2, has a molar mass of 259 g/mol and a solubility of 1.60 g/L at 25 °C.

AB2(s)↽−−⇀A2+(aq)+2B−(aq)

What is the ?sp of this salt at 25 °C?
Answer #1

Solubility = 1.60 gr = 1-60 molla 259 = 0.00610 mol/ AB2 (s) Ataq) + 2 B 609) 0.00618 280.00610 Ksp = [A2+) [B-) Ksp = 60.006